+x 6. Idx X 7. dx 00 -dx 1-COS X
Answer #1

~ 2/5 X + dx 15 +1 1) dx 2/5 -1 지 +1) dx 년 - 3/5 lo +1) dx ~ 3/5 - de + d기 으 -3/5 +! X - + 3/5 11 2/5 5 2/5 개 十C + 2 | 두 2/52. S 7t 7. du 4 2. S (7 7x +2 ) 4 구5 1 2 S (TX+2) af te4 Expandling 2. 13720 3920 50 16807 240 10 32 Ida 기4 29 - 2. 32. 1807M= tank/2 - equin 8 da | - 0082 ܝܐ da = - 1 COXX-1 -] at i-a8²1/2 = soc²X/2 Sec X/2 - 1 | 2 - +1 ser ²1/2 da Q tau212 2 Let I